Kunal is 11 years old and his sisyer medha is 24 years old . After how many years will medha be twice as old as her brother.
Answer by lwsshak3(11628)

Kunal is 11 years old and his sisyer medha is 24 years old . After how many years will medha be twice as old as her brother.
***
let x=years medha will be twice as old as her brother
24+x=2(11+x)
24+x=22+2x
x=2
years medha will be twice as old as her brother=2